OpenStudy (anonymous):

A family spends 20% of its monthly income on food, 23% on rent, and 42% on other expenses and saves the balance. If the family saves $360 per month, what is its monthly income?

OpenStudy (anonymous):

20+23+42=85 So the family saves 15% of their income. So... 360 = income*0.15 income = 360/0.15=$2400

OpenStudy (anonymous):

is it finance..